A company is forecasted to pay dividends of $0.90, $1.20, and $1.45 in 3, 6, and 9 months, respectively. Given interest rates of 5.5%, how much dollar impact will dividends have on option prices? (Assume a 9-month option.)

A) $3.45

B) $3.90

C) $4.22

D) $4.50
Answer #1

Let Dividend after month n be Dn

=> D3 = 0.90
D6 = 1.20
D9 = 1.45

Monthly Interest Rate = r = 0.055/12

Present Value of dividends = D3/(1+r)3 + D6/(1+r)6 + D9/(1+r)9

= 0.90/(1+0.055/12)3 + D6/(1+0.055/12)6 + D9/(1+0.055/12)9

= $3.45